Qd = 100 – 5P and Qs = –20 + 3P. If the price is R10, then there will be a shortage of


1
Expert's answer
2022-04-26T14:54:39-0400

"Q_d=100-5\cdot 10=50,"

"Q_s=-20+3\cdot 10=10,"

"S=Q_d-Q_s=50-10=40."